Grade 11Magical Mathematics[Interesting Approach]

|x^2 - 3x +2| = 2 find x

larisha sharma
13 Years agoGrade 11

1 Answer

ASHOK PAUL
13 Years ago

since the modulus of X^2-3X+2=2 IT COULD BE THAT X^2-3X+2=2 OR X^2-3X+2=-2.But X^2-3X+2=-2 has no real solution,so leave it.

the roots of X^2-3X+2=2 are 3 and 0.so X=3 or 0